OpenStudy (anonymous):

Jennifer is on her daily run. After 25 minutes she has gone 1.5 miles. If she continues at the same pace, how long in minutes will her 8.4 mile run take? I need to make sure I got the right answer I got 0.06 for the rate is that right?

OpenStudy (anonymous):

The rate of 0.06 is correct.

OpenStudy (anonymous):

ok thank you so then how would I begin to figure out how many mins it will take her in 8.4 miles?

OpenStudy (anonymous):

oh wait i think i got it

OpenStudy (anonymous):

I do not know how to find that. I'm sorry. You can search it up or something. I know it's distance over time though to find speed.

OpenStudy (anonymous):

Oh okay, that's good.

OpenStudy (anonymous):

d=r*t

OpenStudy (anonymous):

then you plug in what you know and solve that sounds right to me anyway

OpenStudy (dumbcow):

yes d=r*t 8.4 = .06t 140 = t

OpenStudy (anonymous):

But you don't find the distance in this problem.

OpenStudy (anonymous):

Wait, nevermind.

OpenStudy (anonymous):

wouldn't be 14.0?

OpenStudy (dumbcow):

it cant be 14 because we know she went 1.5 miles in 25 minutes so to go 8.4 miles it would naturally take longer than 25 minutes

OpenStudy (anonymous):

ok thanks